OpenStudy (anonymous):

what is the simplified form of the following expression? (4/7)^3 343/64 21,952 407 64/343

OpenStudy (anonymous):

@Nnesha I'm honestly not sure what this is. Help please? Lol thanks!

Nnesha (nnesha):

here is an example \[\huge\rm (\frac{ a}{ b })^m = \frac{ a^m }{ b^m }\]

OpenStudy (anonymous):

so... \[\frac{ 4^3 }{ 7^3 }\] ?

Nnesha (nnesha):

yes!

OpenStudy (anonymous):

Right here is where I get confused :/

Nnesha (nnesha):

well 4^3 means = 4 times 4 times 4 same with 7 so just multiply them

OpenStudy (anonymous):

64/343 is the answer then?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

Cool. Thank you! :)